OpenStudy (anonymous):

What is the product of three square root of eight times four square root of three? Simplify your answer.

OpenStudy (anonymous):

Please help anybody!

OpenStudy (whpalmer4):

Is the problem \[3\sqrt{8} * 4\sqrt{3}\]?

OpenStudy (whpalmer4):

Remember that multiplication is commutative, so we can rearrange the order without any problem. \[3\sqrt{8}*4\sqrt{3} = 3*4*\sqrt{8}*\sqrt{3} = 12*\sqrt{8}*\sqrt{3}\] One of the properties of square roots is that \[\sqrt{a}*\sqrt{b} = \sqrt{a*b}\] Can you use that property to convert \[12*\sqrt{8}*\sqrt{3}\]into something equivalent but with only one square root sign?